# OfCaM: Global Human Mesh Recovery via Optimization-free Camera Motion Scale Calibration

### v1

@@ 5.3 Comparison with the State-of-the-art
Tab. 3 compares our approach with SOTA world coordinate human mesh recovery methods. Our Method demonstrates a significant enhancement on global human motion metrics over WHAM [28] (about 10% improvement in W-MPJPE, 20% improvement in WA-MPJPE, 60% improvement in RTE). This notable improvement, especially in RTE, which evaluates the entire motion trajectory, is attributed to our accurate camera motion scale calibration. Our calibration effectively and reliably decouples human motion from the camera motion.
@@
Time Complexity. We compare the running time between our method and scale-optimization methods (such as SLAHMR [37] and PACE [14]) in Fig. 1(c). Excluding the SLAM running time, SLAHMR takes over 200 minutes per 1000 frames, and PACE takes 8 minutes per 1000 frames for optimization. In contrast, our approach requires significantly less time (2.5 seconds per 1000 frames) as it is optimization-free. Furthermore, we achieve a better scale compared to SLAHMR, despite the latter’s long optimization process. This illustrates that our scale-calibration is not only time-efficient but also delivers strong performance.
@@ 6 Conclusion & Limitations
Currently, our work has two limitations. First is the body-pose accuracy (see PA-MPJPE error in Table 3). However, our framework is compatible with any HMR model so more advanced methods can be integrated. This is beyond our current scope of accurate recovery of human meshes in world coordinates rather than optimizing local pose metrics. Secondly, like previous work [28], our evaluation of global human and camera motion is limited to the EMDB dataset, as it is the only dataset specifically designed for the global human and camera motion task. Others either lack annotations for world frames (3DPW [34]) or have incomplete data and or code release (HCM [14]).

### v2

@@ 5.3 Comparison with the State-of-the-art
Tab. 3 compares our approach with SOTA world coordinate human mesh recovery methods. Our Method demonstrates a significant enhancement on global human motion metrics over WHAM [WHAM_2023] (about 10% improvement in W-MPJPE, 20% improvement in WA-MPJPE, 60% improvement in RTE). This notable improvement, especially in RTE, which evaluates the entire motion trajectory, is attributed to our accurate camera motion scale calibration. Our calibration effectively and reliably decouples human motion from the camera motion.
@@
Time Complexity. We compare the running time between our method and scale-optimization methods (such as SLAHMR [SLAMHR_2023] and PACE [PACE_2023]) in Fig. 1(c). Excluding the SLAM running time, SLAHMR takes over 200 minutes per 1000 frames, and PACE takes 8 minutes per 1000 frames for optimization. In contrast, our approach requires significantly less time (2.5 seconds per 1000 frames) as it is optimization-free. Furthermore, we achieve a better scale compared to SLAHMR, despite the latter’s long optimization process. This illustrates that our scale-calibration is not only time-efficient but also delivers strong performance.
@@ 6 Conclusion & Limitations
Currently, our work has two limitations. First is the body-pose accuracy (see PA-MPJPE error in Table 3). However, our framework is compatible with any HMR model so more advanced methods can be integrated. This is beyond our current scope of accurate recovery of human meshes in world coordinates rather than optimizing local pose metrics. Secondly, like previous work [WHAM_2023], our evaluation of global human and camera motion is limited to the EMDB dataset, as it is the only dataset specifically designed for the global human and camera motion task. Others either lack annotations for world frames (3DPW [3DPW_2018]) or have incomplete data and or code release (HCM [PACE_2023]).